OpenStudy (anonymous):

How do I write the expression 512^-1/3 in radical notation?

OpenStudy (saifoo.khan):

\[\Huge \sqrt[a]{b} \to b^\frac1a\]

OpenStudy (saifoo.khan):

\[\Huge 512^{-\frac13} »» \frac{1}{512^\frac13} »»\frac{1}{\sqrt[3]{512}}\]

OpenStudy (anonymous):

might help to know that the cubed root of 512 is a nice whole number

OpenStudy (anonymous):

thanks.... I think I just typed in the wrong number into my math lab for my answer. I think I got it now